OpenStudy (anonymous):

In quadrilateral QRST, m∠Q is 68°, m∠R is (3x + 40)°, and m∠T is (5x − 52)°. What are the measures of ∠R , ∠S , and ∠T ? Write the numerical values in that order with the measures separated by commas.

OpenStudy (mathstudent55):

There isn't enough info. Is there a figure?

OpenStudy (anonymous):

it wont let me post it :(

OpenStudy (anonymous):

nvm

OpenStudy (mathstudent55):

Ok. That's the extra info. The quadrilateral is inscribed in a circle.

OpenStudy (mathstudent55):

In an inscribed quadrilateral, opposite angles are supplementary.

OpenStudy (anonymous):

The opposite angles of a quadrilateral inscribed in a circle are supplementary: m∠Q+m∠S= 180° m∠S=180°−m∠Q=180°−68°= 112° m∠R+m∠T=180° (3x + 40) + (5x − 52) = 180° Solving for x, x = 24°. Substituting the value of x in the expressions for m∠R and m∠T , m∠R = 112° and m∠T = 68°.

OpenStudy (campbell_st):

well the opposite angles in a cyclic quadrilateral are supplementary

OpenStudy (mathstudent55):

Angles Q and S are opposite angles. They are supplementary. Can you find the measure of angle S?

OpenStudy (anonymous):

it tells me to put it with commas??? i don't understand ..

OpenStudy (mathstudent55):

1, 2, 3, 4 Numbers separated by commas.

OpenStudy (mathstudent55):

Only 3 numbers, though.

OpenStudy (anonymous):

ok ooohhh i get it ok :) thank you! ^.^

OpenStudy (mathstudent55):

yw

OpenStudy (anonymous):

so how would it actually be with the commas?

OpenStudy (anonymous):

can someone help me?

OpenStudy (anonymous):

m∠Q+m∠S= 180° m∠S=180°−m∠Q=180°−68°= 112° m∠R+m∠T=180° (3x + 40) + (5x − 52) = 180° Solving for x, x = 24°. Substituting the value of x in the expressions for m∠R and m∠T , m∠R = 112° and m∠T = 68°

OpenStudy (anonymous):

how do i write the answer

OpenStudy (anonymous):

112,112,68 can i get a medal :)